OpenStudy (anonymous):

How do I do this? 3x - 5 <10?

OpenStudy (dangerousjesse):

Substitute :3 Do you know how to do that?

OpenStudy (anonymous):

All i could do was this 3x - 5 x 10 +5 +5

OpenStudy (anonymous):

Not really ;/

OpenStudy (f_jayyy):

switch 5 to the to the other side and make it a positive

OpenStudy (anonymous):

so 3x +5 <10?

OpenStudy (f_jayyy):

no 3x < 10 + 5 then 3x < 15 then divide both sides by three and you get x < 5

OpenStudy (kendricklamar2014):

You need to add 5 to both sides: \[3x−5+5<10+5\] @AnimeWover What is 10 + 5?

OpenStudy (anonymous):

10+ 5 = 15

OpenStudy (kendricklamar2014):

Good

OpenStudy (f_jayyy):

then divide both sides by three and you get ???? x < 5

OpenStudy (anonymous):

What does the X mean? Why is it there though?

OpenStudy (dangerousjesse):

Okay, that's fine :) \[\large3x-5<10\]\[\large-5+5=0\]\[\large10+5=15\]\[\large3x<15\] As 3x is less than fifteen, not equal to, \[\large x\neq5\]

OpenStudy (constanceleak):

you are trying to find the unknown number

OpenStudy (constanceleak):

x represents the unknown number

OpenStudy (f_jayyy):

x is the variable all it is the num ber you are trying to find

OpenStudy (kendricklamar2014):

Now, we divide both sides by 3: \[\frac{ 3x }{ 3 } < \frac{ 15 }{ 3 }\] \[\frac{ 3x }{ 3 } = x\] WHat is 15/3 @AnimeWover

OpenStudy (anonymous):

15 / 3 = 5

OpenStudy (f_jayyy):

soooo x is 5

OpenStudy (dangerousjesse):

X cannot equal 5.

OpenStudy (dangerousjesse):

3x is less than 15, not 15

OpenStudy (constanceleak):

x<5 right?

OpenStudy (dangerousjesse):

Therefore x is less than 5, not 5.

OpenStudy (f_jayyy):

ik that if u were to look at the equation i did above that what i had...

OpenStudy (kendricklamar2014):

x < 5 is the answer

OpenStudy (f_jayyy):

i had that all the way on the top lol

OpenStudy (kendricklamar2014):

Do you understand? @AnimeWover

OpenStudy (f_jayyy):

just one rule to remember when you switch a number to either side of the equal sign ALWAYS change the sign..if its positive make it a negative and if its a negative make it a positive

OpenStudy (kendricklamar2014):

Yes ^

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so 3x + 5 < 10 5+5 10+5

OpenStudy (dangerousjesse):

Huh? I thought you made it zero and added/subtracted it from the sum/product/etc.

OpenStudy (anonymous):

3 / 15 = 5 x < 5

OpenStudy (anonymous):

I think i'm skipping steps. Sorry

OpenStudy (anonymous):

i am a lost cause with math XD

OpenStudy (kendricklamar2014):

Step 1: Add 5 to both sides. Step 2: Divide both sides by 3

OpenStudy (dangerousjesse):

Yes, that makes more sense^

OpenStudy (f_jayyy):

so it ends up being 3x - 5 = 10 3x = 10 + 5 3x = 15 3/3 x = 15/3 x = 5

OpenStudy (dangerousjesse):

F_Jay, we do not know what x is equal to. We only know it is less than 5.

OpenStudy (f_jayyy):

lol sorry i keep forgtrtting im so used to the = sign

OpenStudy (anonymous):

Wait hoiw are we diving BOTH sides?

OpenStudy (f_jayyy):

to get x by itself

OpenStudy (anonymous):

oh

OpenStudy (dangerousjesse):

No problem :) We are dividing both sides by 3 because x is being multiplied, and the end result is the product of x.

OpenStudy (kendricklamar2014):

I'll put it more in detail: Step 1: Add 5 to both sides: \[3x−5+5<10+5\] The 5's the the left side cancel out since: -5 + 5 = 0. On the right side, 10 + 5 = 15. \[3x<15\] Step 2: Divide both sides by 3: \[\frac{ 3x }{ 3 } < \frac{ 15 }{ 3 }\] On the left side, 3 divided by 3 = 1x, 1x is the same as x. On the right side, 15 divided by 3 = 5. \[Answer: x < 5\]

OpenStudy (f_jayyy):

yup :)

OpenStudy (anonymous):

ohhh

OpenStudy (kendricklamar2014):

Yup, does that help you understand?

OpenStudy (anonymous):

how do we know X is less than 5 though?

OpenStudy (kendricklamar2014):

because: < = Less than > = Greater than

OpenStudy (f_jayyy):

this is how i remember those signs.... < faces the negative side ( negative=less than) and > positive=greater than) faces the positive side

OpenStudy (anonymous):

ok, makes much more sense XD

OpenStudy (kendricklamar2014):

\[:)\]